OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5x-25/5 5x − 5 x − 25 x + 5 x − 5

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

\[\frac{5x - 25}{5} = \frac{5(x - 5)}{5} = \frac{\cancel{5}(x - 5)}{\cancel{5}} = x - 5\]